# Testing Low-Resource Language Support in LLMs Using Language Proficiency Exams: the Case of Luxembourgish

### v2

@@ Abstract
Large Language Models (LLMs) have become an increasingly important tool in research and society at large. While LLMs are regularly used all over the world by experts and lay-people alike, they are predominantly developed with English-speaking users in mind, performing well in English and other wide-spread languages while less-resourced languages such as Luxembourgish are seen as a lower priority. This lack of attention is also reflected in the sparsity of available evaluation tools and datasets. In this study, we investigate the viability of language proficiency exams as such evaluation tools for the Luxembourgish language. We find that large models such as ChatGPT, Claude and DeepSeek-R1 typically achieve high scores, while smaller models show weak performances. We also find that the performances in such language exams can be used to predict performances in other NLP tasks.
@@ 1 Introduction
As such, it is important to determine the capabilities of models to understand and generate texts in such languages, and to recognise when they can fail. The benefits of such research are two-fold as it allows (1) to find general weaknesses and pitfalls that can be addressed and mitigated in future LLMs; (2) to make recommendations on which LLM to use for a given scenario in a given low-resource language. The latter is especially important in use cases where an LLM would have to handle sensitive data such as financial or medical information that is not allowed to leave a financial or medical institution due to internal policies and data privacy laws such as the Health Insurance Portability and Accountability Act (HIPAA) Act (1996) or the the General Data Protection Regulation (GDPR) Regulation (2016). In such cases, data must be processed using locally deployed models that do not access the internet.
In this study, we aim to investigate the capabilities of LLMs for Luxembourgish, a low-resource language spoken by nearly 600 000 people world-wide111according to https://cursus.edu/en/23040/luxembourgish-at-its-best, but predominantly in Luxembourg. We want to determine those capabilities from two angles, (a) how well LLMs perform in linguistic proficiency exams used in language institutions; (b) whether their performance in such exams can predict their performance in text generation tasks.
Specifically, we investigate 45 LLMs from a wide array of LLM families that we compare systematically. In addition, we ascertain in which cases LLMs fail and discover trends and common weaknesses using language exams. Finally, we aim to determine whether or not such exams are a useful tool to predict how well an LLM performs in two tasks related to text summarisation. While this study focuses on the Luxembourgish language, we believe that our methodology can be generalised to other low-resource languages as well.
@@
    A systematic comparison of 45 popular LLMs with regard to their linguistic proficiency in Luxembourgish
@@
    A quantitative and qualitative error analysis
@@
    An analysis on the usefulness of language exams as a tool for performance prediction on text generation tasks.
@@
    An accessible public leaderboard that can be used to recommend LLMs for given use cases [^2] The leaderboard can be found at https://applux-nnjdvksvbpgovzux7dtqts.streamlit.app/
@@ 2 Methodology
In this section, we give details on the methodology of this study and on our experiments, including the datasets we used, the LLMs we targeted, the metrics to compare the LLMs, and the prompt design. We also present and elaborate on the research questions we address in this paper.
We evaluate LLMs in two general disciplines: (1) language fluency and (2) language generation. The language fluency test consists of solving language exams that comprise sets of multiple-choice questions that relate to the general understanding of the language itself. In order to evaluate language generation, we use two tasks from the LuxGen benchmark Plum et al. (2025), specifically the headline generation and short description tasks. These tasks both consist of processing texts and generating appropriate summaries, either in form of a short description or a news headline.
Figure 1 shows the general testing pipeline. We first initiate the model with a unified prompt across every LLM (1). This initiation prompt consists of explaining the LLM’s role and the general task it needs to solve, as well as specifying the format of the inputs and outputs. We then pre-process the test samples to make them uniform and compatible with our prompts (2), and give them to the LLM in random order (3). For the language exams, we apply a post-processing step before comparing it to the ground truth (4). Post-processing generally consists of correcting minor typos in case the LLM’s output is different from each of the provided answers (however, we judge outputs as being invalid if they deviate significantly or are ambiguous and could be attributed to multiple answers). Finally, the LLM’s performance will be measured using task-specific metrics (see Section 2.4).
Figure 1: General Testing Pipeline
@@ 2.1 Research Questions
  3. RQ3: Is there a correlation between performing well in a language exam and performing well in a language generation task? For the final question, we determine whether using language exams is an appropriate method to decide an LLM’s language capabilities and if the "winners" of the exams manage to perform well in generation tasks.
@@
Total 104 104 102 112 105 103 630
Table 1: Statistics for language tests
2.2 Datasets
2.2.1 Language Tests
In order to evaluate the fluency of each model, we use official language exams for Luxembourgish provided to us by the Institut National des Langues Luxembourg, a language institute in Luxembourg [^3] https://www.inll.lu/en/. The difficulty level of each test follows the Common European Framework of Reference for Languages (CEFR) guidelines, and thusly ranges from A1 (beginner) to C2 (proficient) [^4] Further explanations can be found at https://www.efset.org/cefr/. The tests consist of mixtures of multiple-choice questions (MCQs) and open fill-in-the-blank type questions. However, for our experiments, we unify the exams by transforming them into the same type of questions. One of the authors who speaks the language at a native level converted the open questions into MCQs as this allows to use a single standardised prompt across all tests, and simplifies the error analysis. Questions fall into one of four broad categories: vocabulary, grammar, reading comprehension (RC), and listening comprehension. Note that the data for the latter category consists of transcripts rather than audio files. As such, we will refer to that category as conversation comprehension (CC) in this paper. Table 1 shows further details for the tests.
@@ 2.2.3 Short Description Generation
Small Models ($\leq 15B$) Medium-sized Models ($>15B\&\leq 200B$)
Qwen 2 0.5 26.9 22.1 23.3 22.3 19.0 18.4 Mistral-Small 22 4.8 3.8 2.9 0.0 1.9 1.0
Llama 3.2 1 1.9 3.8 1.0 4.5 11.4 3.9 Gemma 2 27 21.2 6.7 3.9 1.8 2.9 2.9
Qwen 2 1.5 32.7 27.9 26.2 32.1 30.5 22.3 Aya-23 35 7.7 3.8 1.0 3.6 4.8 1.0
StableLM 2 1.6 0.0 14.4 0.0 17.0 9.5 15.5 Command-R 35 6.7 1.9 1.0 0.9 1.9 0.0
Gemma 2 2 39.4 32.7 30.1 0.0 36.2 28.2 Alfred 40 1.0 0.0 1.0 0.0 3.8 1.0
Llama 3.2 3 37.5 35.6 28.2 36.6 31.4 23.3 Mixtral 8x7 1.0 0.0 0.0 0.0 0.0 1.0
Phi 3 3.8 45.2 36.5 27.2 26.8 26.7 17.5 Llama 2 70 1.9 0.0 0.0 0.0 0.0 1.0
Phi 3.5 3.8 6.7 12.5 1.9 1.8 16.2 21.4 Llama 3 70 17.3 11.5 9.7 9.8 10.5 1.9
Qwen 2 7 48.1 36.5 34.0 35.7 36.2 33.0 Llama 3.1 70 21.2 12.5 33.0 10.7 11.4 1.9
Llama 2 7 7.7 10.6 8.7 13.4 23.8 20.4 DeepSeek-R1 70 17.3 13.4 9.7 7.1 8.6 7.8
WizardLM 2 7 24.0 30.8 27.2 19.6 35.2 23.3 Qwen 2 72 19.2 9.6 2.9 4.5 1.9 2.9
Mistral 7 35.6 35.6 35.0 31.3 31.4 26.2 Command-R+ 104 13.5 9.6 2.9 1.8 5.7 1.0
Aya-23 8 40.4 30.8 29.1 31.3 28.6 33.0 Mistral-Large 123 16.3 5.8 6.8 5.4 7.6 3.9
Llama 3.1 8 56.7 33.7 33.0 33.9 40.0 33.0 Mixtral 8x22 2.9 0.0 0.0 0.0 1.9 1.9
Llama 3 8 45.2 35.6 35.0 35.7 31.4 31.1 WizardLM 2 8x22 9.6 3.8 3.9 0.9 1.9 1.9
DeepSeek-R1 8 42.3 31.7 35.0 41.1 26.7 26.2 Large Models ($>200B$)
GLM 4 9 54.8 41.3 38.8 45.5 38.1 34.0 Llama 3.1 405 78.8 39.4 49.5 59.8 56.2 47.6
Gemma 2 9 58.7 47.1 39.8 0.0 37.1 33.0 DeepSeek-R1 671 92.3 83.7 74.8 76.8 72.4 82.5
Mistral-Nemo 12 52.9 36.5 30.1 34.8 28.6 21.4 ChatGPT 3.5 unk. 92.3 85.6 81.6 83.9 77.1 65.0
StableLM 2 12 27.9 7.7 26.2 29.5 25.7 24.3 Claude 3.5 Sonnet unk. 92.3 92.3 81.6 88.4 77.1 68.9
Llama 2 13 27.9 0.0 0.0 0.0 22.9 23.3 ChatGPT 4o unk. 92.3 85.6 79.6 84.8 74.3 60.2
Phi 3 14 45.2 38.5 39.8 39.3 35.2 24.3 ChatGPT 4o mini unk. 90.4 90.4 60.2 59.8 55.2 43.7
Gemini 2.0 Flash unk. 89.4 84.6 75.7 81.3 75.2 54.4
LeChat unk. 42.3 53.8 39.8 52.7 41.9 32.0
@@ 2.3 Models
We perform our experiments on a large number of LLMs from a wide array of language model families and model sizes. The main criteria for including a given LLM in our selection is their reported ability to (a) understand Luxembourgish, (b) understand a closely related Germanic language such as German or Dutch, or (c) understand a very large number of languages even if Luxembourgish is not explicitly included. Specifically, we include the following models: Alfred Hallström et al. (2023), Aya-23 Aryabumi et al. (2024), ChatGPT (OpenAI, 2023, OpenAI, 2024a, OpenAI, 2024b), Claude 3.5 Sonnet Anthropic (2024), Command-R Cohere For AI (2024), DeepSeek-R1 DeepSeek-AI (2025), Gemini 2.0 Flash Google , LeChat Mistral AI (2023a), Gemma 2 Team (2024), GLM 4 GLM et al. (2024), Llama 2 Touvron et al. (2023), Llama 3, Llama 3.1, Llama 3.2 Grattafiori et al. (2024), Mistral Mistral AI (2023b), Mixtral Mistral AI (2024), Phi 3, Phi 3.5 Abdin et al. (2024), Qwen 2 qwe (2024), StableLM 2 Bellagente et al. (2024), and WizardLM 2 Xu et al. (2024). Appendix B shows the full list of LLMs together with model sizes and reasons for inclusion.
Model sizes range from 1 billion parameters (Llama3.2:1B) to 671 billion (DeepSeek-R1:671B). We divide them into 3 size categories: small ($\leq 15B$ parameters), medium-sized ($>15B$& $\leq 200B$ parameters), and large ($>200B$ parameters). Note that although some model sizes are unknown to us as they are not communicated by the companies releasing them, they are estimated to be at least hundreds of billions of parameters. As such, we categorise the concerned LLMs as large.
@@ 2.4 Metrics
For language generation, we evaluate the outputs of LLMs in terms of grammar and spelling as well as similarity to the ground truth. We measure the performance on both of these tasks in an LLM-as-a-judge fashion using an LLM that is not part of the ones we are investigating as the judging LLM. We choose the judging LLM based on its ability to solve the task at hand, which we manually verify. The metrics used should meet the following requirements:
@@
      range between 00 and $1$ where 00 is incomprehensible and $1$ is perfect with neither spelling nor grammar mistakes.
@@
      punish non-Luxembourgish outputs, i.e., return 00 if the output is in a language other than Luxembourgish.
  • Ground Truth Similarity metric
@@
      range between 00 and $1$ where $1$ is identical to the ground truth and decreases the less similar the output is to the ground truth and the more it deviates from the given text.
@@
Considering that such LLM-generated metrics are essentially black boxes and non-deterministic due to the inherent random nature of LLMs, we repeat the generation a total of three times in order to reduce variations.

### v4

@@ Abstract
Large Language Models (LLMs) have become an increasingly important tool in research and society at large. While LLMs are regularly used all over the world by experts and lay-people alike, they are predominantly developed with English-speaking users in mind, performing well in English and other wide-spread languages while less-resourced languages such as Luxembourgish are seen as a lower priority. This lack of attention is also reflected in the sparsity of available evaluation tools and datasets. In this study, we investigate the viability of language proficiency exams as such evaluation tools for the Luxembourgish language. We find that large models such as Claude and DeepSeek-R1 typically achieve high scores, while smaller models show weak performances. We also find that the performances in such language exams can be used to predict performances in other NLP tasks in Luxembourgish.
@@ 1 Introduction
As such, it is important to determine the capabilities of models to understand and generate texts in such languages, and to recognise when they can fail. The benefits of such research are two-fold as it allows (1) to find general weaknesses and pitfalls that can be addressed and mitigated in future LLMs; (2) to make recommendations on which LLM to use for a given scenario in a given low-resource language. The latter is especially important in use cases where an LLM would have to handle personally identifiable data such as personal names or phone numbers that is not allowed to leave a company or organisation due to internal policies and data privacy laws such as the Luxembourgish Data Protection Law Grand-Duché de Luxembourg (2018) or the General Data Protection Regulation European Union (2016). In such cases, data must be processed using locally deployed models that do not access the internet.
In this study, we aim to investigate the capabilities of LLMs for Luxembourgish, a low-resource language spoken by nearly 600,000 people world-wide [^1] https://cursus.edu/en/23040/luxembourgish-at-its-best, but predominantly in Luxembourg. We want to determine those capabilities from two angles, (a) how well LLMs perform in linguistic proficiency exams used in language institutions; (b) whether their performance in such exams can predict their performance in Natural Language Generation (NLG) tasks.
Specifically, we investigate 53 LLMs from a wide array of LLM families that we compare systematically. In addition, we ascertain in which cases LLMs fail and discover trends and common weaknesses using language exams. Finally, we aim to determine whether or not such exams are a useful tool to predict how well an LLM performs in two tasks related to text summarisation. While this study focuses on the Luxembourgish language, we believe that our methodology can be generalised to other low-resource languages as well.
@@
    A systematic comparison of 53 popular LLMs with regard to their linguistic proficiency in Luxembourgish.
@@
    A quantitative and qualitative error analysis.
@@
    An analysis on the usefulness of language exams as a tool for performance prediction on NLG tasks for Luxembourgish.
@@
    An accessible and interactive public leaderboard for performance in Luxembourgish that we will extend as new LLMs are released. [^2] The leaderboard can be found as part of our AI Sandbox at https://ai-sandbox.list.lu/
2 Related Work
2.1 Testing LLMs on Knowledge Exams
2.2 Testing Linguistic capabilities of LLMs with Language exams
However, compared to related studies, we do not only assess LLM performance on language exams, but we also explore their usefulness to predict LLM performance in general NLP tasks. Furthermore, we use tests that follow the Common European Framework of Reference for Languages (CEFR) Council of Europe (2001) to evaluate LLMs.
@@ 3 Methodology
In this section, we give details on the methodology of this study and on our experiments, including the datasets we used, the LLMs we targeted, the metrics to compare the LLMs, the prompt design, and hyperparameters of the studied models. We also present and elaborate on the research questions we address in this paper.
We evaluate LLMs in two general disciplines: (1) language fluency and (2) language generation. The language fluency test consists of solving language exams that comprise sets of multiple-choice questions that relate to the general understanding of the language itself. In order to evaluate NLG, we use two tasks from the LuxGen benchmark Plum et al. (2025), specifically the Headline Generation and Short Description tasks. These tasks both consist of processing texts and generating appropriate summaries, either in form of a short description or a news headline.
Testing is done in a zero-shot fashion for all tasks. Figure 1 shows the general testing pipeline. After instantiating an LLM (1), for every test sample, we initiate the model with a prompt that is identical across every LLM (2). This initiation prompt consists of explaining the LLM’s role and the general task it needs to solve, as well as specifying the format of the inputs and outputs. We then pre-process the test samples to make them uniform and compatible with our prompts (3), and give them to the LLM one at a time and in random order (4). For the language exams, we apply a post-processing step before comparing it to the ground truth (5). Post-processing generally consists of automatically correcting minor typos if the output is different from each of the provided answers (however, we judge outputs as being invalid if they deviate significantly or are ambiguous and could be attributed to multiple answers). Finally, the LLM’s performance will be measured using task-specific metrics (6) (see Section 3.4).
Figure 1: General testing pipeline.
@@ 3.1 Research Questions
  3. RQ3: Is there a correlation between performing well in a language exam and performing well in an NLG task? For the final question, we determine whether using language exams is an appropriate method to decide an LLM’s language capabilities and if the "winners" of the exams manage to perform well in generation tasks.
3.2 Datasets
3.2.1 Language Tests
@@
Total 104 104 102 114 104 101 629
Table 1: Statistics for language tests.
In order to evaluate the fluency of each model, we use official language exams for Luxembourgish provided to us by the Institut National des Langues Luxembourg (INLL), a language institute in Luxembourg [^3] https://www.inll.lu/en/. The difficulty level of each test follows CEFR guidelines, and thusly ranges from A1 (beginner) to C2 (proficient). The tests consist of fill-in-the-blank type questions. The large majority of them are multiple-choice questions (MCQs) while nearly 17% of them them were open without a pre-determined selection of possible answers. The number of open questions differs depending on the CEFR level and are either related to grammar or vocabulary. Appendix A shows the exact number and distribution of open questions in the exams. The A1 and A2 levels do not contain any open questions, then, from B1 level, the higher the level the higher the number of open questions.
However, for our experiments, we unify the exams by transforming them into the same type of questions. Two Luxembourgish-speaking authors converted the open questions into MCQs as this allows to use a single standardised prompt across all tests, and simplifies the error analysis. The designated authors made sure that the difficulty of the converted questions is adequate for the level of the corresponding exams by writing distractors that matched the difficulty of the given exam, taking other MCQ questions from the same level as inspiration. Distractors for grammar-related questions include lexically similar misspellings of the correct answer, incorrectly conjugated verbs, and incorrect usage of prepositions. Vocabulary-related distractors generally include lexically similar words to the correct answer and near-synonyms to the correct answer that are wrong in the context of the given text. Questions fall into one of four broad categories: vocabulary, grammar, reading comprehension (RC), and listening comprehension. Note that the data for the latter category consists of transcripts rather than audio files. As such, we will refer to that category as conversation comprehension (CC) in this paper. Table 6 shows further details for the tests. Further note that we do not have permission to give access to the actual dataset, but we can establish contact between interested researchers and the INLL.
@@ 3.2.3 Short Description Generation
Baseline (average chance rate) Baseline (average chance rate)
Random Baseline / 37.6 37.5 33.3 33.0 31.8 33.5 Random Baseline / 37.6 37.5 33.3 33.0 31.8 33.5
[…]
Qwen 2 0.5 26.9 22.1 23.3 22.8 22.1 15.8 Mistral-Small 22 51.9 46.2 31.1 39.5 38.5 25.7
Llama 3.2 1 26.0 19.2 28.2 21.1 19.2 12.9 Gemma 2 27 66.3 63.5 54.4 45.6 50.0 41.6
Gemma 3 1 45.2 36.5 38.8 40.4 25.0 33.7 Gemma 3 27 71.2 72.1 62.1 60.5 60.6 44.6
Qwen 2 1.5 32.7 27.9 25.2 31.6 21.2 23.8 QWQ 32 60.6 49.0 51.5 47.4 38.5 44.6
StableLM 2 1.6 22.1 14.4 21.4 16.7 15.4 16.8 Aya-23 35 43.3 37.5 39.8 36.0 33.7 34.7
EuroLLM 1.7 17.3 21.2 12.6 23.7 26.0 17.8 Command-R 35 51.0 38.5 33.0 42.1 35.6 37.6
Gemma 2 2 39.4 32.7 31.1 32.5 27.9 28.7 Alfred 40 42.3 39.4 29.1 29.8 36.5 29.7
Llama 3.2 3 37.5 35.6 26.2 35.1 30.8 24.8 Mixtral 8x7 50.0 41.3 34.0 39.5 31.7 29.7
Phi 3 3.8 45.2 36.5 27.2 26.3 27.9 17.8 Llama 2 70 38.5 33.7 28.2 39.5 26.0 39.6
Phi 3.5 3.8 39.4 27.9 11.7 10.5 7.7 10.9 Llama 3 70 52.9 49.0 38.8 35.1 38.5 43.6
Gemma 3 4 54.8 43.3 39.8 46.5 36.5 31.7 Llama 3.1 70 55.8 56.7 49.5 52.6 47.1 41.6
Qwen 2 7 48.1 36.5 32.0 36.0 39.4 31.7 DeepSeek-R1 70 64.4 51.0 56.3 54.4 49.0 41.6
Llama 2 7 36.5 38.5 24.3 39.5 25.0 26.7 Qwen 2 72 60.6 46.2 40.8 51.8 41.3 40.6
WizardLM 2 7 44.2 38.5 21.4 31.6 28.8 29.7 Command-R+ 104 52.9 38.5 39.8 44.7 31.7 32.7
Mistral 7 35.6 35.6 34.0 30.7 34.6 26.7 Mistral-Large 123 66.3 50.0 42.7 52.6 43.3 44.6
Aya-23 8 40.4 30.8 30.1 31.6 31.7 31.7 Mixtral 8x22 62.5 49.0 41.7 42.1 39.4 39.6
Llama 3.1 8 56.7 33.7 33.0 33.3 30.8 33.7 WizardLM 2 8x22 58.7 40.4 40.8 43.0 35.6 34.7
Llama 3 8 45.2 35.6 35.0 36.0 28.8 31.7 Large Models (>200>200B)
DeepSeek-R1 8 53.8 39.4 38.8 39.5 34.6 28.7 Llama 3.1 405 78.8 39.4 48.5 59.6 55.8 46.5
GLM 4 9 54.8 41.3 37.9 45.6 44.2 35.6 DeepSeek-R1 671 92.3 83.7 73.8 76.3 73.1 81.2
EuroLLM 9 44.2 45.2 33.0 34.2 31.7 25.7 ChatGPT 3.5 unk. 92.3 85.6 80.6 83.3 76.0 66.3
Gemma 2 9 58.7 47.1 38.8 49.1 40.4 34.7 Claude 3.5 Sonnet unk. 92.3 92.3 80.6 86.8 77.9 69.3
Mistral-Nemo 12 52.9 36.5 28.2 33.3 29.8 20.8 ChatGPT 4o unk. 92.3 85.6 78.6 84.2 74.0 60.4
StableLM 2 12 27.9 35.6 26.2 28.9 27.9 23.8 ChatGPT 4o mini unk. 90.4 90.4 59.2 59.6 54.8 43.6
Gemma 3 12 62.5 58.7 53.4 50.9 47.1 40.6 Gemini 2.0 Flash unk. 89.4 84.6 74.8 80.7 76.0 54.5
Llama 2 13 27.9 14.4 23.3 26.3 23.1 26.7 LeChat unk. 42.3 53.8 38.8 51.8 41.3 32.7
Phi 3 14 45.2 38.5 38.8 39.5 27.9 21.8
Phi 4 14 53.8 51.9 42.7 44.7 34.6 33.7
@@ 3.3 Models
We perform our experiments on a large number of LLMs from a wide array of language model families and model sizes. As models are not explicitly reported to understand Luxembourgish, the main criteria for including a given LLM in our selection is their reported ability to (a) understand a closely related Germanic language such as German or Dutch, or (b) understand a very large number of languages. Specifically, we include the following models: Alfred Hallström et al. (2023), Aya-23 Aryabumi et al. (2024), ChatGPT (OpenAI, 2023, OpenAI, 2024a, OpenAI, 2024b), Claude 3.5 Sonnet Anthropic (2024), Command-R Cohere For AI (2024), DeepSeek-R1 DeepSeek-AI (2025), EuroLLM Martins et al. (2024), Gemini 2.0 Flash Google (2024), LeChat Mistral AI (2023a), Gemma 2 Mesnard et al. (2024), GLM 4 Zeng et al. (2024), Gemma 3 Kamath et al. (2025), Llama 2 Touvron et al. (2023), Llama 3, Llama 3.1, Llama 3.2 Grattafiori et al. (2024), Mistral Mistral AI (2023b), Mixtral Mistral AI (2024), Phi 3, Phi 3.5 Abdin et al. (2024a), Phi 4 Abdin et al. (2024b), Qwen 2 Yang et al. (2024a), QWQ (Yang et al., 2024b, Qwen, 2025), StableLM 2 Bellagente et al. (2024), and WizardLM 2 Xu et al. (2024). Appendix B shows the full list of LLMs together with model sizes and reasons for inclusion.
Model sizes range from 0.5 billion parameters (Qwen2:0.5B) to 671 billion (DeepSeek-R1:671B). We divide them into 3 size categories: small ($\leq 15$B parameters), medium-sized (>15>15B & $\leq 200$B parameters), and large (>200>200B parameters). Note that although some model sizes are unknown to us as they are not communicated by the companies releasing them, they are estimated to be at least hundreds of billions of parameters. As such, we categorise the concerned LLMs as large.
@@ 3.4 Metrics
For language generation, we evaluate the outputs of LLMs in terms of grammar and spelling as well as similarity to the ground truth. We measure the performance on both of these tasks in an LLM-as-a-judge (LaaJ) fashion using an LLM that is not part of the ones we are investigating as the judging LLM. We choose the judging LLM based on its ability to solve the task at hand, which we manually verify. The metrics should meet the following requirements:
@@
      range between 0 and $1$ where 0 is incomprehensible and $1$ is perfect with neither spelling nor grammar mistakes.
@@
      punish non-Luxembourgish outputs, i.e., return 0 if the output is written in a different language.
  • Adequacy metric
@@
      range between 0 and $1$ where $1$ is identical to the ground truth and decreases the less similar the output is to the ground truth and the more it deviates from the text on a semantic level.
@@
Considering that such LLM-generated metrics are essentially black boxes and non-deterministic due to the inherent random nature of LLMs, we repeat the generation a total of three times in order to reduce variation. In addition to the LaaJ metrics, we use two traditional metrics: METEOR Banerjee and Lavie (2005) and BERTScore Zhang et al. (2019), using the Luxembourgish LuxemBERT Lothritz et al. (2022) as the embedding model for BERTScore.